in 2015 a company estimated that 65% of its customers were females n ,in that year,3 customers entered 1of the shop's. Assuming that these customers were not related or acquaintances,find the probability that they were
1)All females
2)All male
3)At least one male
1
Expert's answer
2020-03-23T14:02:01-0400

Let "X=" the number of customers who are females: "X\sim B(n;p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "p=0.65, n=3"

1) The probability that these customers were all females is


"P(X=3)=\binom{3}{3}0.65^3(1-0.65)^{3-3}=0.274625"

2) The probability that these customers were all males is


"P(X=0)=\binom{3}{0}0.65^0(1-0.65)^{3-0}=0.042875"

3) The probability that at least one male is


"P(X<3)=1-P(X=3)=1-0.274625=0.725375"